The slope whose equation is 3x-4y=7
Answer by nerdybill(7384)

The slope whose equation is 3x-4y=7
.
put given equation into "slope-intercept" form
y = mx + b
where m is the slope
.
3x-4y=7
-4y=-3x+7
4y=3x-7
y=(3/4)x-(7/4)
.
Slope, then, from inspection is
3/4